Find two consecutive even integers such that the sum of the larger and twice the smaller is 62.
Answer by Lenzky(18)

Let x be the smaller nunber and x+2 be the larger
x+2 + 2x = 62
combine similar terms
x+2x = 62-2
3x=60 divide both side by 3 to get x
x=20
x+2=?
20+2= 22
20 & 22
CHECKING
22+ (2)20 = 62
22+40=62
62 = 62